How many grams of ice at -19.2 ∘C can be completely converted to liquid at 25.3 ∘C if the available heat for this process is 4.57×103 kJ ? For ice, use a specific heat of 2.01 J/(g⋅∘C) and ΔHfus=6.01kJ/mol .
Answer #1

hos, -19.2°C] net ho [1, 25.30] 1120 [s, ovc] - H200,00] Onet = Q + d + Oz let mass of wle = wg. Q = mSoT - W .x2.01x (19.2)answer may vary a little as I have taken specific heat of liquid water to be 4.184jouke per gram